OpenStudy (anonymous):

solve(x+2)^2=1 please help best answer will get the medal. Please helppppp!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

(x+2)^2=1 (x+2)(x+2)=1 x^2+4x+4=1 x^2+4x=3 4x=3-x^2 x=(3-x^2)\4

OpenStudy (anonymous):

The right answer is =-3 =-1 but i need to know how to solve it.

jigglypuff314 (jigglypuff314):

(x + 2)^2 = (x + 2)(x + 2) = x^2 + 4x + 4 so you get x^2 + 4x + 4 = 1 - 1 -1 x^2 + 4x + 3 = 0 then factor so (x + 3)(x + 1) = 0 so solve for x when x + 3 = 0 and x + 1 = 0 that would give you your answers :)

OpenStudy (anonymous):

Hi, First you need to "get rid" of that exponent by distributing it to the parenthesis. So, you need to: 1. Square the first term x^2 2. Multiply the two terms and double them 4x 3. square the last term 4 Making your final answer for what is in parenthesis : x^2 + 4x + 4 Set this equal to 1 (it's in the problem that way) then subtract it, leaving the final equation as: x^2 +4x + 3 = 0 Now factor: (x+3)(x+1)= 0 and solve for x! Hope that helps :)

OpenStudy (anonymous):

Glad to help!

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Any time!

OpenStudy (anonymous):

could you help with another Q

OpenStudy (anonymous):

Sure!